A Complete Bibliography of Publications in Journal of Computational Physics: X

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 December 2023 Version 1.00

Title word cross-reference

2 [AE21]. 3 [BSD^{+}20, FDB23, NGP^{+}19].

2019 [Ano19e, Ano19f, Ano19g, Ano19h]. 2020 [Ano20e, Ano20f, Ano20g, Ano20h]. 2021 [Ano21e, Ano21f, Ano21g, Ano21h]. 2022 [Ano22e, Ano22f, Ano22g, Ano22h]. 2023 [Ano23b].

acceleration [Chu23]. Active [vLtI19, WA23]. Adaptive [CCFH20, HVH20, AVPK19, CACM19, FNSA19, KHAA19, MCF^{+}21, PMWK20]. ADM [CCFH20]. ADM-LTS [CCFH20]. ALE [MTA23]. ALE-ISPH [MTA23]. algorithm [HGL^{+}19, TR19]. algorithms [Tso19]. alloy [PMWK20]. ambiguity [ESJ22]. amplifiers [NGP^{+}19]. analysis [RKvdM21, VLB^{+}19]. anisotropic [Chu23]. Application [DGJT19, AE21, CHMP23, HOL19, WL19]. applications [HGL^{+}19, vLtI19]. applied [KMP^{+}20]. approach [CBS20, DDC23, NGP^{+}19, WJNH19]. approaches [VOdLL19]. aqueous [SPNSB19]. arbitrarily [DKKEB20]. arbitrary [EDvW20a, EDvW20b, TR19]. arising [LAH22]. arithmetic [FAA20]. arithmetic-average [FAA20]. artificial [BHL22]. artificial-viscosity [BHL22]. assignment [BHL22]. average [FAA20]. axisymmetric [WPIJ20].

Back [WL19]. based [BHL22, LCK19, LMN^{+}20, MTA23, MCF^{+}21]. bases [AHS^{+}19]. basis [Chu23]. Bayesian [BE22]. binary [PMWK20]. blocks [Chu23]. Board [Ano19a, Ano19b, Ano19c, Ano19d, Ano20a, Ano20b, Ano20c, Ano20d, Ano21a, Ano21b, Ano21c, Ano21d, Ano22a, Ano22b, Ano22c, Ano22d, Ano23a]. bodies [DKKEB20]. Boltzmann [LYH19]. bottom [FP20]. boundaries [HOL19]. boundary [HR20, MTA23]. bounded [CP23]. brain [RMPA20]. bubbly [KMP^{+}20].

Cahn [CWWW19]. calculus [LVB^{+}19, VLB^{+}19]. capture [KG20]. capturing [ZP23]. cardiac [Chu23]. Carlo [PRB23, SP19]. Cartesian [EDvW20b, VOdLL19]. Cartesian-grid [VOdLL19]. Cell [FDB23, GKHG23, LMN^{+}20, MCF^{+}21, FDB22]. centred [HGL^{+}19]. charged [WPIJ20]. chemical [SPNSB19]. Comparisons [VOdLL19]. compensation [WL19]. compressible [CP23, CFC^{+}20, VOdLL19, WWGS21]. Computational [BSD^{+}20, WA23]. Computing [LCK19]. concurrent [RKvdM21]. condition [MTA23]. conditional [PDF19]. conditions [HR20]. conduction [Chu23]. conductivity [DGJT19]. conformal [DDC23]. conformally [FNSA19]. Connecting [SPH^{+}20]. Consecutive [Mar21]. conservation [DGJT19, KHAA19]. conservative [UkM19]. considerations [HOL19]. constitutive [RKvdM21]. construction [RKvdM21]. continuous [VCSSNP19]. continuum [CBS20]. convection [WWGS21]. convergence [XP21]. coordinates [LVB^{+}19, VLB^{+}19]. Corrected [MTA23, GORV21]. correction [WL19]. counterpart [DGJT19]. coupling [SP19, WJNH19]. crystals [MMR20, VCSSNP19]. Cubic [Mar21]. curvature [EDvW20b]. curved [Chu23]. cut [AE21].

D [FDB23, NGP^{+}19, AE21, BSD^{+}20]. data [BKTZ19, BE22]. deep [BE22]. defects [MMR20]. deformable [HOL19]. dense [KMP^{+}20]. density [MTA23, SPH^{+}20]. density-based [MTA23]. derivations [VLB^{+}19]. derivatives [AHS^{+}19]. design [VCSSNP19]. detectors [PRB23]. deterministic [DGJT19]. difference [LLR21]. differential [BE22]. differentiation [Dub19]. dilute [EDvW20a]. dimensional [Del19, DDC23, PDF19, WPIJ20]. direct [JLZ19, SP19]. Discontinuous [ZP23, CACM19, CFC^{+}20, GKL^{+}19, TGC19]. discovery [BE22]. discrete [WWGS21]. discretization [HR20, LLR21]. dissipations [UkM19]. Divergence [FNSA19]. Divergence-free [FNSA19]. DNS [CP23]. domains [ABS19, HR20]. DPG [NGP^{+}19]. drift [HSvL20]. driven [CHMP23]. dynamic [HGL^{+}19, HVH20]. dynamics [BSD^{+}20, BHL22, LCK19, SPH^{+}20, WA23]. dynamos [CDRT20].

eddy [GKL^{+}19]. Editorial [Abg19, Abg21, Ano19a, Ano19b, Ano19c, Ano19d, Ano20a, Ano20b, Ano20c, Ano20d, Ano21a, Ano21b, Ano21c, Ano21d, Ano22a, Ano22b, Ano22c, Ano22d, Ano23a]. EEG [RMPA20]. Efficient [BKTZ19, GKHG23, WJNH19]. electric [Chu23, WPIJ20]. electromagnetic [BKTZ19]. element [AE21, AVPK19, GKL^{+}19, XP21, ZP23]. Elliptical [FDB23, FDB22]. Energy [HOL19, AVPK19, CWWW19, SPH^{+}20]. energy-density [SPH^{+}20]. ENO [HM21]. Ensemble [RBMM23a, RBMM23b]. EPIC [FDB23, FDB22]. equal [HSvL20]. equal-weights [HSvL20]. equation [BE22, CHWW20, CWWW19, Del19, HR20]. equations [CACM19, RMPA20, WL19]. error [WL19]. escape [KG20]. estimate [FAA20]. estimation [EDvW20b]. Euler [EDvW20a, KMP^{+}20, LÖR23]. Eulerian [FP20]. evaluation [GORV21]. examples [LVB^{+}19]. expansion [LÖR23]. explicit [GKL^{+}19, HGL^{+}19]. extended [PRB23]. extrapolation [Dub19].

factor [PBSM23]. family [RMPA20]. Fast [GORV21, PBSM23, HGL^{+}19, KG20, SP19]. FC [BHL22]. FC-based [BHL22]. fiber [NGP^{+}19]. fibers [RMPA20]. field [AVPK19, BKTZ19, LÖR23, WPIJ20]. fields [DGJT19, TR19]. films [LCK19]. filter [HSvL20]. filtering [GKL^{+}19]. finite [AE21, AVPK19, HGL^{+}19, LLR21, VOdLL19]. flat [FP20]. flow [CBS20, DDC23, LMN^{+}20, VOdLL19, WJNH19]. flows [AE21, CHMP23, CP23, EDvW20a, KMP^{+}20, WWGS21]. fluid [BSD^{+}20, Mar21, XP21]. fluid-structure [XP21]. fluids [DKKEB20, WA23]. flux [vLtI19]. fly [RKvdM21]. forecasting [HSvL20]. formulae [Dub19]. formulations [HGL^{+}19]. Fort [Del19]. forth [WL19]. forward [RMPA20]. fractured [HVH20]. framework [RBMM23a, VOdLL19]. Frankel [Del19]. free [AE21, CHMP23, FNSA19, LLR21, LAH22, WJNH19]. free-surface [LAH22]. front [UkM19]. front-tracking [UkM19]. FSI [HOL19]. fully [AVPK19]. function [EDvW20b]. functions [AHS^{+}19].

gain [NGP^{+}19]. Galerkin [CACM19, CFC^{+}20, TGC19, VCSSNP19, ZP23]. gas [LYH19, WWGS21]. general [Mar21]. geometric [Chu23, vLtI19]. geometry [GKHG23]. geothermal [HVH20]. glacier [AE21]. GPU [LCK19, SP19]. GPU-based [LCK19]. grid [CP23, MCF^{+}21, TGC19, VOdLL19]. grid-based [MCF^{+}21]. grids [EDvW20b].

heat [LYH19, SP19]. Height [EDvW20b]. Height-function [EDvW20b]. Heisenberg [LÖR23]. Heterogeneous [SPH^{+}20, ABS19, CCFH20, DGJT19]. Hierarchical [ESJ22, PBSM23]. High [FP20, CACM19, HM21, SPH^{+}20, VOdLL19]. High-Order [FP20, VOdLL19]. high-resolution [HM21]. highly [DKKEB20]. Hilliard [CWWW19]. HLL [FAA20]. HLL-type [FAA20]. homogeneous [CHWW20]. hp} [CACM19]. hp}-adaptive [CACM19]. Hybrid [HM21, LYH19]. hybridized [CFC^{+}20]. Hydrodynamics [DKKEB20]. hyperbolic [KHAA19, PDF19, WL19].

ice [LAH22]. ice-sheet [LAH22]. II [LVB^{+}19, RBMM23b]. Implementation [LVB^{+}19, SP19]. Implicit [CFC^{+}20, HSvL20, LMN^{+}20]. improved [WWGS21]. incompressible [CACM19, HGL^{+}19, VOdLL19]. inconsistency [FAA20]. Increasing [LAH22]. information [BKTZ19]. inhomogeneous [TR19]. integral [RMPA20]. integration [CDRT20]. interaction [SPNSB19, XP21]. interface [Mar21]. interpolation [Mar21]. invariants [Del19]. inverse [ESJ22, JLZ19]. irradiance [PBSM23]. Isospectral [ABS19]. ISPH [MTA23]. Iterative [Mar21].

Jacobi [LVB^{+}19, VLB^{+}19]. January [Ano19e, Ano20e, Ano21e, Ano22e]. June [Ano19f, Ano20f, Ano21f, Ano22f].

kinematic [CDRT20]. kinetic [SPH^{+}20, WWGS21].

laden [EDvW20a]. Lagrange [EDvW20a, KMP^{+}20]. Lagrangian [FP20, HGL^{+}19, SPNSB19]. Landau [CHWW20]. Laplace [HR20]. large [GKL^{+}19, KMP^{+}20, LCK19]. laser [CHMP23, NGP^{+}19]. lattice [LYH19]. laws [DGJT19, KHAA19]. layer [GORV21]. leapfrog [WPIJ20]. learning [BE22, RKvdM21]. linear [AVPK19, WL19]. liquid [MMR20]. local [KHAA19, LMN^{+}20]. localized [BHL22]. locally [GORV21]. locally-corrected [GORV21]. logarithmic [CWWW19]. low [PBSM23]. low-rank [PBSM23]. LTS [CCFH20].

machine [RKvdM21]. magnetic [TR19, WPIJ20]. magnetohydrodynamics [CFC^{+}20, FNSA19]. mapping [DDC23]. March [Ano19g, Ano20g, Ano21g, Ano22g]. Massively [HSvL20]. Mathematical [VLB^{+}19]. matrices [PBSM23]. matter [SPH^{+}20]. Maxwell [NGP^{+}19, TGC19, WL19]. mechanical [RKvdM21]. mechanism [PRB23]. media [CCFH20, CBS20, LMN^{+}20, WJNH19]. melting [CHMP23]. mesh [EDvW20a, MMR20, PMWK20]. mesh-spacing [EDvW20a]. meshes [FNSA19, Tso19, vLtI19]. method [AE21, AVPK19, CHWW20, CACM19, FP20, FNSA19, FDB22, FDB23, GKL^{+}19, HOL19, HM21, LYH19, LLR21, LMN^{+}20, MMR20, PDF19, SPNSB19, SPH^{+}20, UkM19, VCSSNP19, WL19, ZP23]. methodology [XP21]. methods [CFC^{+}20, GORV21, JLZ19, LLR21, TGC19]. micro [CBS20]. micro-continuum [CBS20]. missing [UkM19]. mixed [FP20].

Modal [GKL^{+}19]. model [KMP^{+}20, WA23]. modeling [CBS20, RMPA20, WJNH19]. Modelling [PMWK20, AE21, AVPK19, DDC23, EDvW20a, MMR20]. models [RKvdM21]. molecular [SPH^{+}20]. moments [PDF19]. Monte [SP19, PRB23]. moving [FNSA19, MMR20, vLtI19]. multilevel [CCFH20, HVH20]. Multiphase [CBS20, LMN^{+}20]. multipole [GORV21]. multiresolution [KHAA19]. multiscale [CBS20, RKvdM21, SPH^{+}20, VCSSNP19]. multiwavelet [AHS^{+}19].

narrow [KG20]. Natural [CP23, WWGS21]. Navier [CACM19]. nearly [HGL^{+}19]. nematic [MMR20]. network [BHL22, WJNH19]. Neumann [HR20, MTA23]. neural [BHL22]. neural-network [BHL22]. Newtonian [AE21]. noise [MCF^{+}21]. noisy [BE22]. non [AE21, EDvW20b]. non-Newtonian [AE21]. non-uniform [EDvW20b]. Nonlinear [RBMM23b, LÖR23, NGP^{+}19]. novel [MTA23]. November [Ano23b]. Numerical [CHMP23, LÖR23, ABS19, CWWW19, Dub19, SP19, XP21]. Numerically [UkM19].

ocean [FP20, HSvL20]. off [PRB23]. On-the-fly [RKvdM21]. one [Del19]. one-dimensional [Del19]. open [CBS20]. open-source [CBS20]. optics [vLtI19]. optimization [ESJ22, GKHG23]. Order [FP20, AVPK19, CACM19, EDvW20b, TR19, VOdLL19]. organized [WA23]. orthonormal [Chu23].

Parallel [CDRT20, CACM19, HSvL20]. Parallel-in-time [CDRT20]. Parallelization [KMP^{+}20]. parameter [BE22]. Parcel [FDB23, FDB22]. Parcel-In-Cell [FDB23, FDB22]. Part [LVB^{+}19, RBMM23a, RBMM23b, VLB^{+}19]. partial [BE22]. Particle [DKKEB20, CHWW20, EDvW20a, HSvL20, MTA23, MCF^{+}21]. particle-in-cell [MCF^{+}21]. particle-laden [EDvW20a]. particle-size [EDvW20a]. particles [TR19, WPIJ20]. parts [ZP23]. peel [PRB23]. Phase [JLZ19, AVPK19, BKTZ19, SPNSB19]. phase-field [AVPK19]. Phaseless [JLZ19]. phenomena [AVPK19]. photonic [VCSSNP19]. physical [UkM19]. planetary [PBSM23]. point [BSD^{+}20]. polishing [CHMP23]. polygonal [HR20]. polyhedrons [Mar21]. polynomials [LVB^{+}19, VLB^{+}19]. pore [WJNH19]. pore-network [WJNH19]. porous [CCFH20, CBS20, LMN^{+}20, WJNH19]. positioning [Mar21]. Positivity [CWWW19]. Positivity-preserving [CWWW19]. potential [CWWW19, FNSA19]. potentials [GORV21]. preserving [CWWW19]. probabilistic [RKvdM21]. problem [JLZ19, LAH22, RMPA20]. problems [DGJT19, ESJ22, HOL19, KG20]. process [Dub19]. propagation [Chu23, FP20]. properties [Del19, ZP23].

quadrature [PDF19]. quadratures [GORV21]. quantification [DGJT19]. quantum [LÖR23].

radiative [PRB23, SP19]. Raman [NGP^{+}19]. random [LYH19]. rank [PBSM23]. ratio [EDvW20a]. RBF [HM21]. RBF-ENO [HM21]. reactive [SPNSB19]. real [BKTZ19]. real-valued [BKTZ19]. recursive [LLR21]. reduction [MCF^{+}21]. refinement [PMWK20]. regularization [ESJ22]. Relative [Chu23]. Relativistic [Del19, WPIJ20]. remark [Dub19]. reordering [LMN^{+}20]. represented [AHS^{+}19]. reservoirs [HVH20]. resolution [HM21]. retrieval [BKTZ19, JLZ19]. Richardson [Dub19]. Riemann [FAA20]. rigid [DKKEB20]. robust [VCSSNP19].

sampling [JLZ19]. scale [KMP^{+}20, LCK19]. scattering [JLZ19]. scheme [CCFH20, Del19, DKKEB20, KHAA19, WWGS21, WPIJ20]. schemes [CWWW19, MCF^{+}21, Tso19, vLtI19]. Schrödinger [Del19]. second [AVPK19]. second-order [AVPK19]. selection [Tso19]. self [WA23]. self-organized [WA23]. September [Ano19h, Ano20h, Ano21h, Ano22h]. set [BSD^{+}20]. shape [GKHG23, WA23]. shaped [DKKEB20]. sheet [LAH22]. shifting [MTA23]. shock [BHL22, ZP23]. shock-dynamics [BHL22]. signal [Chu23, FAA20]. simulating [UkM19, WWGS21]. simulation [CHMP23, GKL^{+}19, HVH20, SP19, VCSSNP19]. simulations [KHAA19, LCK19, LÖR23, LAH22]. single [LLR21]. single-step [LLR21]. size [EDvW20a]. sizes [LAH22]. smooth [AHS^{+}19]. Smoothed [DKKEB20]. smoothing [RBMM23a, RBMM23b]. solid [HGL^{+}19, SPNSB19]. solid/aqueous [SPNSB19]. solidification [PMWK20]. solids [LYH19]. solution [ESJ22]. solver [BHL22, KG20]. solvers [FAA20]. source [CBS20, JLZ19]. space [CCFH20]. space-time-stepping [CCFH20]. spacing [EDvW20a]. Sparse [MCF^{+}21, TGC19, BE22]. spatially [PRB23]. Spectral [FP20, GKL^{+}19, XP21, ZP23]. spectral-element [XP21]. speed [FAA20]. SPH [HOL19]. sphere [KG20]. spherical [LVB^{+}19, VLB^{+}19]. Spline [Mar21]. stable [AVPK19, CWWW19, LAH22]. static [WPIJ20]. Stencil [Tso19]. step [LLR21, LAH22]. stepping [CCFH20, KHAA19, LMN^{+}20, TR19]. stochastic [DGJT19, KMP^{+}20, VCSSNP19]. Stokes [CACM19]. strategy [MCF^{+}21]. stratified [DDC23]. stretching [CP23]. structure [XP21]. study [ABS19, XP21]. summation [ZP23]. surface [AE21, CHMP23, FP20, LAH22]. surfaces [BSD^{+}20, Chu23, PBSM23, WA23]. surrogate [RKvdM21]. system [LLR21, TGC19]. system-free [LLR21]. systems [LYH19, WL19].

technique [MTA23]. temperature [CHMP23]. temperature-driven [CHMP23]. temporal [LLR21]. Tensor [LVB^{+}19, VLB^{+}19]. theory [SPH^{+}20]. thermal [PBSM23]. thin [LCK19]. Three [PDF19]. Three-dimensional [PDF19]. time [CCFH20, CDRT20, KHAA19, LMN^{+}20, LAH22, TR19]. time-step [LAH22]. time-stepping [KHAA19, LMN^{+}20, TR19]. Total [HGL^{+}19]. tracking [TR19, UkM19]. trajectories [WPIJ20]. trajectory [HSvL20]. transfer [LYH19, PRB23, SP19]. transport [CCFH20, RBMM23a, RBMM23b, SPNSB19]. truly [HGL^{+}19, VOdLL19]. Two [WPIJ20, DDC23]. Two-dimensional [WPIJ20, DDC23]. type [FAA20].

uncertainty [DGJT19]. underdetermined [ESJ22]. Unified [RBMM23a, WWGS21]. uniform [EDvW20b]. uniqueness [JLZ19]. unstructured [Tso19]. Updated [HGL^{+}19]. updates [RBMM23b]. upwind [HGL^{+}19]. using [FNSA19, LVB^{+}19, VLB^{+}19, WJNH19].

vacuum [LÖR23]. valued [BKTZ19]. variable [GKHG23]. vector [FNSA19]. vectors [Chu23]. Verification [XP21]. vertex [HGL^{+}19]. via [LCK19, RMPA20]. view [PBSM23]. viscosity [BHL22]. viscous [DKKEB20, VOdLL19]. Vlasov [TGC19]. volume [HGL^{+}19, Mar21, VOdLL19]. volume-of-fluid [Mar21].

walk [LYH19]. wall [CP23]. wall-bounded [CP23]. waves [FP20]. weak [LÖR23]. weak-field [LÖR23]. weakly [VOdLL19]. weakly-compressible [VOdLL19]. weights [HSvL20]. WENO [Tso19]. wetting [AVPK19]. wire [RMPA20]. within [DKKEB20].

References

Abgrall:2019:E

[Abg19] Rémi Abgrall. Editorial. Journal of Computational Physics: X, 3:??, June 2019. CODEN JCPXAL. ISSN 2590-0552.

            URL http://www.sciencedirect.com/science/article/pii/
            S2590055219300423.

Abgrall:2021:E

[Abg21] Rémi Abgrall. Editorial. Journal of Computational Physics: X, 11:??, June 2021. CODEN JCPXAL. ISSN 2590-0552.

            URL http://www.sciencedirect.com/science/article/pii/
            S2590055221000172.

Amore:2019:IHD

[ABS19] Paolo Amore, John P. Boyd, and Natalia Tene Sandoval. Isospectral heterogeneous domains: a numerical study. Journal of Computational Physics: X, 1:??, January 2019. CODEN

            JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
            com/science/article/pii/S2590055219300344.

Ahlkrona:2021:CFE

[AE21] Josefin Ahlkrona and Daniel Elfverson. A cut finite element method for non-Newtonian free surface flows in 2D — application to glacier modelling. Journal of Computational Physics: X, 11:??, June 2021. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S259005522100007X.

Anderson:2019:DSF

[AHS^{+}19] Joel Anderson, Robert J. Harrison, Hideo Sekino, Bryan Sundahl, Gregory Beylkin, George I. Fann, Stig R. Jensen, and Irina Sagert. On derivatives of smooth functions represented in multiwavelet bases. Journal of Computational Physics: X, 4:??, September 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300496.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board. Journal of Computational Physics: X, 1:i, January 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300435.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. Journal of Computational Physics: X, 2:??, March 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300502.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. Journal of Computational Physics: X, 3:??, June 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300563.

Anonymous:2019:EBd

[Ano19d] Anonymous. Editorial Board. Journal of Computational Physics: X, 4:??, September 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300605.

Anonymous:2019:Ja

[Ano19e] Anonymous. January 2019. Journal of Computational Physics: X, 1:??, January 2019. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2019:Jb

[Ano19f] Anonymous. June 2019. Journal of Computational Physics: X, 3: ??, June 2019. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2019:M

[Ano19g] Anonymous. March 2019. Journal of Computational Physics: X, 2:??, March 2019. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2019:S

[Ano19h] Anonymous. September 2019. Journal of Computational Physics: X, 4:??, September 2019. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2020:EBa

[Ano20a] Anonymous. Editorial Board. Journal of Computational Physics: X, 5:??, January 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300068.

Anonymous:2020:EBb

[Ano20b] Anonymous. Editorial Board. Journal of Computational Physics: X, 6:??, March 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300159.

Anonymous:2020:EBc

[Ano20c] Anonymous. Editorial Board. Journal of Computational Physics: X, 7:??, June 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300263.

Anonymous:2020:EBd

[Ano20d] Anonymous. Editorial Board. Journal of Computational Physics: X, 8:??, September 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300329.

Anonymous:2020:Ja

[Ano20e] Anonymous. January 2020. Journal of Computational Physics: X, 5:??, January 2020. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2020:Jb

[Ano20f] Anonymous. June 2020. Journal of Computational Physics: X, 7: ??, June 2020. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2020:M

[Ano20g] Anonymous. March 2020. Journal of Computational Physics: X, 6:??, March 2020. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2020:S

[Ano20h] Anonymous. September 2020. Journal of Computational Physics: X, 8:??, September 2020. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2021:EBa

[Ano21a] Anonymous. Editorial Board. Journal of Computational Physics: X, 9:??, January 2021. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055221000020.

Anonymous:2021:EBb

[Ano21b] Anonymous. Editorial Board. Journal of Computational Physics: X, 10:??, March 2021. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055221000123.

Anonymous:2021:EBc

[Ano21c] Anonymous. Editorial Board. Journal of Computational Physics: X, 11:??, June 2021. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055221000184.

Anonymous:2021:EBd

[Ano21d] Anonymous. Editorial Board. Journal of Computational Physics: X, 12:??, September 2021. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055221000202.

Anonymous:2021:Ja

[Ano21e] Anonymous. January 2021. Journal of Computational Physics: X, 9:??, January 2021. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2021:Jb

[Ano21f] Anonymous. June 2021. Journal of Computational Physics: X, 11:??, June 2021. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2021:M

[Ano21g] Anonymous. March 2021. Journal of Computational Physics: X, 10:??, March 2021. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2021:S

[Ano21h] Anonymous. September 2021. Journal of Computational Physics: X, 12:??, September 2021. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2022:EBa

[Ano22a] Anonymous. Editorial Board. Journal of Computational Physics: X, 13:??, January 2022. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055222000026.

Anonymous:2022:EBb

[Ano22b] Anonymous. Editorial Board. Journal of Computational Physics: X, 14:??, March 2022. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055222000075.

Anonymous:2022:EBc

[Ano22c] Anonymous. Editorial Board. Journal of Computational Physics: X, 15:??, June 2022. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055222000129.

Anonymous:2022:EBd

[Ano22d] Anonymous. Editorial Board. Journal of Computational Physics: X, 16:??, September 2022. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055222000178.

Anonymous:2022:Ja

[Ano22e] Anonymous. January 2022. Journal of Computational Physics: X, 13:??, January 2022. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2022:Jb

[Ano22f] Anonymous. June 2022. Journal of Computational Physics: X, 15:??, June 2022. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2022:M

[Ano22g] Anonymous. March 2022. Journal of Computational Physics: X, 14:??, March 2022. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2022:S

[Ano22h] Anonymous. September 2022. Journal of Computational Physics: X, 16:??, September 2022. CODEN JCPXAL. ISSN 2590-0552.

Anonymous:2023:EB

[Ano23a] Anonymous. Editorial Board. Journal of Computational Physics: X, 17:??, November 2023. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S259005522300015X.

Anonymous:2023:N

[Ano23b] Anonymous. November 2023. Journal of Computational Physics: X, 17:??, November 2023. CODEN JCPXAL. ISSN 2590-0552.

Aymard:2019:LSO

[AVPK19] Benjamin Aymard, Urbain Vaes, Marc Pradas, and Serafim Kalliadasis. A linear, second-order, energy stable, fully adaptive finite element method for phase-field modelling of wetting phenomena. Journal of Computational Physics: X, 2:??, March 2019. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055219300265.

Bonneville:2022:BDL

[BE22] Christophe Bonneville and Christopher Earls. Bayesian deep learning for partial differential equation parameter discovery with sparse and noisy data. Journal of Computational Physics: X, 16:??, September 2022. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055222000117.

Bruno:2022:FBS

[BHL22] Oscar P. Bruno, Jan S. Hesthaven, and Daniel V. Leibovici. FC-based shock-dynamics solver with neural-network localized artificial-viscosity assignment. Journal of Computational Physics: X, 15:??, June 2022. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055222000063.

Blinne:2019:ERP

[BKTZ19] Alexander Blinne, Stephan Kuschel, Stefan Tietze, and Matt Zepf. Efficient retrieval of phase information from real-valued electromagnetic field data. Journal of Computational Physics: X, 1:??, January 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300356.

Bouchiba:2020:CFD

[BSD^{+}20] Hassan Bouchiba, Simon Santoso, Jean-Emmanuel Deschaud, Luisa Rocha-Da-Silva, François Goulette, and Thierry Coupez. Computational fluid dynamics on 3D point set surfaces. Journal of Computational Physics: X, 7:??, June 2020. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055220300214.

Chalmers:2019:PHA

[CACM19] N. Chalmers, G. Agbaglah, M. Chrust, and C. Mavriplis. A parallel hp-adaptive high order discontinuous Galerkin method for the incompressible Navier–Stokes equations. Journal of Computational Physics: X, 2:??, March 2019. CODEN JCPXAL. ISSN

2590-0552. URL http://www.sciencedirect.com/science/
article/pii/S2590055219300393.

Carrillo:2020:MFM

[CBS20] Francisco J. Carrillo, Ian C. Bourg, and Cyprien Soulaine. Multiphase flow modeling in multiscale porous media: an open-source micro-continuum approach. Journal of Computational Physics: X, 8:??, September 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300251.

Carciopolo:2020:AMS

[CCFH20] Ludovica Delpopolo Carciopolo, Matteo Cusini, Luca Formaggia, and Hadi Hajibeygi. Adaptive multilevel space-time-stepping scheme for transport in heterogeneous porous media (ADM-LTS). Journal of Computational Physics: X, 6:??, March 2020. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055220300044.

Clarke:2020:PTI

[CDRT20] Andrew T. Clarke, Christopher J. Davies, Daniel Ruprecht, and Steven M. Tobias. Parallel-in-time integration of kinematic dynamos. Journal of Computational Physics: X, 7:??, June 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300093.

Ciuca:2020:IHD

[CFC^{+}20] C. Ciucă, P. Fernandez, A. Christophe, N. C. Nguyen, and J. Peraire. Implicit hybridized discontinuous Galerkin methods for compressible magnetohydrodynamics. Journal of Computational Physics: X, 5:??, January 2020. CODEN JCPXAL. ISSN

2590-0552. URL http://www.sciencedirect.com/science/
article/pii/S2590055219300587.

Caboussat:2023:NST

[CHMP23] Alexandre Caboussat, Julien Hess, Alexandre Masserey, and Marco Picasso. Numerical simulation of temperature-driven free surface flows, with application to laser melting and polishing. Journal of Computational Physics: X, 17:??, November 2023. CODEN JCPXAL. ISSN 2590-0552. URL http://www.

sciencedirect.com/science/article/pii/S2590055223000057.

Chun:2023:RAO

[Chu23] Sehun Chun. Relative acceleration of orthonormal basis vectors for the geometric conduction blocks of the cardiac electric signal propagation on anisotropic curved surfaces. Journal of Computational Physics: X, 17:??, November 2023. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055223000136.

Carrillo:2020:PMH

[CHWW20] Jose A. Carrillo, Jingwei Hu, Li Wang, and Jeremy Wu. A particle method for the homogeneous Landau equation. Journal of Computational Physics: X, 7:??, June 2020. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055220300184.

Ceci:2023:NGS

[CP23] Alessandro Ceci and Sergio Pirozzoli. Natural grid stretching for DNS of compressible wall-bounded flows. Journal of Computational Physics: X, 17:??, November 2023. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055223000069.

Chen:2019:PPE

[CWWW19] Wenbin Chen, Cheng Wang, Xiaoming Wang, and Steven M. Wise. Positivity-preserving, energy stable numerical schemes for the Cahn–Hilliard equation with logarithmic potential. Journal of Computational Physics: X, 3:??, June 2019. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055219300472.

Dritschel:2023:CMA

[DDC23] Heidi J. Dritschel, David G. Dritschel, and Magda Carr. A conformal mapping approach to modelling two-dimensional stratified flow. Journal of Computational Physics: X, 17: ??, November 2023. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055223000070.

Dellar:2019:RPI

[Del19] Paul J. Dellar. Relativistic properties and invariants of the Du Fort–Frankel scheme for the one-dimensional Schrödinger equation. Journal of Computational Physics: X, 2:??, March 2019. CODEN JCPXAL. ISSN 2590-0552. URL http://www.

sciencedirect.com/science/article/pii/S2590055219300034.

Delgoshaie:2019:SCC

[DGJT19] Amir H. Delgoshaie, Peter W. Glynn, Patrick Jenny, and Hamdi A. Tchelepi. The stochastic counterpart of conservation laws with heterogeneous conductivity fields: Application to deterministic problems and uncertainty quantification. Journal of Computational Physics: X, 2:??, March 2019. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055219300046.

Dietemann:2020:SPH

[DKKEB20] Bastien Dietemann, Torsten Kraft, Harald Kruggel-Emden, and Claas Bierwisch. A Smoothed Particle Hydrodynamics scheme for arbitrarily shaped rigid bodies within highly viscous fluids. Journal of Computational Physics: X, 8:??, September 2020. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055220300202.

Dubeau:2019:RRE

[Dub19] François Dubeau. A remark on Richardson’s extrapolation process and numerical differentiation formulae. Journal of Computational Physics: X, 2:??, March 2019. CODEN JCPXAL. ISSN

2590-0552. URL http://www.sciencedirect.com/science/
article/pii/S2590055219300332.

Evrard:2020:ELM

[EDvW20a] Fabien Evrard, Fabian Denner, and Berend van Wachem. Euler– Lagrange modelling of dilute particle-laden flows with arbitrary particle-size to mesh-spacing ratio. Journal of Computational Physics: X, 8:??, September 2020. CODEN JCPXAL. ISSN

2590-0552. URL http://www.sciencedirect.com/science/
article/pii/S2590055220300305.

Evrard:2020:HFC

[EDvW20b] Fabien Evrard, Fabian Denner, and Berend van Wachem. Heightfunction curvature estimation with arbitrary order on nonuniform Cartesian grids. Journal of Computational Physics: X, 7:??, June 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300123.

Epp:2022:HRS

[ESJ22] Robert Epp, Franca Schmid, and Patrick Jenny. Hierarchical regularization of solution ambiguity in underdetermined inverse and optimization problems. Journal of Computational Physics: X, 13:??, January 2022. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055222000014.

Fleischmann:2020:IAA

[FAA20] Nico Fleischmann, Stefan Adami, and Nikolaus A. Adams. On an inconsistency of the arithmetic-average signal speed estimate for HLL-type Riemann solvers. Journal of Computational Physics: X, 8:??, September 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300299.

Frey:2022:EEP

[FDB22] Matthias Frey, David Dritschel, and Steven Böing. EPIC: the Elliptical Parcel-In-Cell method. Journal of Computational Physics: X, 14:??, March 2022. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055222000051.

Frey:2023:EPC

[FDB23] Matthias Frey, David Dritschel, and Steven Böing. The 3D Elliptical Parcel-In-Cell (EPIC) method. Journal of Computational Physics: X, 17:??, November 2023. CODEN JCPXAL. ISSN

2590-0552. URL http://www.sciencedirect.com/science/
article/pii/S2590055223000148.

Fragile:2019:DFM

[FNSA19] P. Chris Fragile, Daniel Nemergut, Payden L. Shaw, and Peter Anninos. Divergence-free magnetohydrodynamics on conformally moving, adaptive meshes using a vector potential method. Journal of Computational Physics: X, 2:??, March 2019. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055219300368.

Fouques:2020:MEL

[FP20] Sébastien Fouques and Csaba Pákozdi. A mixed Eulerian– Lagrangian high-order spectral method for the propagation of ocean surface waves over a flat bottom. Journal of Computational Physics: X, 8:??, September 2020. CODEN JCPXAL. ISSN

2590-0552. URL http://www.sciencedirect.com/science/
article/pii/S2590055220300238.

Gubler:2023:EVC

[GKHG23] Moritz Gubler, Marco Krummenacher, Hannes Huber, and Stefan Goedecker. Efficient variable cell shape geometry optimization. Journal of Computational Physics: X, 17: ??, November 2023. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055223000094.

Ghiasi:2019:MEF

[GKL^{+}19] Zia Ghiasi, Jonathan Komperda, Dongru Li, Ahmad Peyvan, David Nicholls, and Farzad Mashayek. Modal explicit filtering for large eddy simulation in discontinuous spectral element method. Journal of Computational Physics: X, 3:??, June 2019. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S259005521930040X.

Greengard:2021:FMM

[GORV21] Leslie Greengard, Michael O’Neil, Manas Rachh, and Felipe Vico. Fast multipole methods for the evaluation of layer potentials with locally-corrected quadratures. Journal of Computational Physics: X, 10:??, March 2021. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055221000093.

Hassan:2019:UVC

[HGL^{+}19] Osama I. Hassan, Ataollah Ghavamian, Chun Hean Lee, Antonio J. Gil, Javier Bonet, and Ferdinando Auricchio. An upwind vertex centred finite volume algorithm for nearly and truly incompressible explicit fast solid dynamic applications: Total and updated Lagrangian formulations. Journal of Computational Physics: X, 3:??, June 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300411.

Hesthaven:2021:HHR

[HM21] Jan S. Hesthaven and Fabian Mönkeberg. Hybrid high-resolution RBF-ENO method. Journal of Computational Physics: X, 12: ??, September 2021. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055221000068.

Hermange:2019:ECS

[HOL19] C. Hermange, G. Oger, and D. Le Touzé. Energy considerations in the SPH method with deformable boundaries and application to FSI problems. Journal of Computational Physics: X, 1:??, January 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300241.

Hoskins:2020:DLE

[HR20] Jeremy Hoskins and Manas Rachh. On the discretization of Laplace’s equation with Neumann boundary conditions on polygonal domains. Journal of Computational Physics: X, 8: ??, September 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S259005522030024X.

Holm:2020:MPI

[HSvL20] Håvard Heitlo Holm, Martin Lilleeng Sætra, and Peter Jan van Leeuwen. Massively parallel implicit equal-weights particle filter for ocean drift trajectory forecasting. Journal of Computational Physics: X, 6:??, March 2020. CODEN JCPXAL. ISSN

2590-0552. URL http://www.sciencedirect.com/science/
article/pii/S2590055220300056.

HosseiniMehr:2020:ADM

[HVH20] Mousa HosseiniMehr, Cornelis Vuik, and Hadi Hajibeygi. Adaptive dynamic multilevel simulation of fractured geothermal reservoirs. Journal of Computational Physics: X, 7:??, June 2020. CODEN JCPXAL. ISSN 2590-0552. URL http://www.

sciencedirect.com/science/article/pii/S2590055220300135.

Ji:2019:PIS

[JLZ19] Xia Ji, Xiaodong Liu, and Bo Zhang. Phaseless inverse source scattering problem: Phase retrieval, uniqueness and direct sampling methods. Journal of Computational Physics: X, 1:??, January 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300022.

Kaye:2020:FSN

[KG20] Jason Kaye and Leslie Greengard. A fast solver for the narrow capture and narrow escape problems in the sphere. Journal of Computational Physics: X, 5:??, January 2020. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055219300630.

Kaiser:2019:ALT

[KHAA19] Jakob W. J. Kaiser, Nils Hoppe, Stefan Adami, and Nikolaus A. Adams. An adaptive local time-stepping scheme for multiresolution simulations of hyperbolic conservation laws. Journal of Computational Physics: X, 4:??, September 2019. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S259005521930054X.

Kamath:2020:PSE

[KMP^{+}20] S. Kamath, M. V. Masterov, J. T. Padding, K. A. Buist, M. W. Baltussen, and J. A. M. Kuipers. Parallelization of a stochastic Euler–Lagrange model applied to large scale dense bubbly flows. Journal of Computational Physics: X, 8:??, September 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S259005522030010X.

Lofgren:2022:IST

[LAH22] André Löfgren, Josefin Ahlkrona, and Christian Helanow. Increasing stable time-step sizes of the free-surface problem arising in ice-sheet simulations. Journal of Computational Physics: X, 16:??, September 2022. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055222000105.

Lam:2019:CDT

[LCK19] Michael-Angelo Y.-H. Lam, Linda J. Cummings, and Lou Kondic. Computing dynamics of thin films via large scale GPU-based simulations. Journal of Computational Physics: X, 2:??, March 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055218300015.

Lee:2021:RSF

[LLR21] Youngjun Lee, Dongwook Lee, and Adam Reyes. A recursive system-free single-step temporal discretization method for finite difference methods. Journal of Computational Physics: X, 12:??, September 2021. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055221000159.

Linga:2020:ILT

[LMN^{+}20] Gaute Linga, Olav Møyner, Halvor Møll Nilsen, Arthur Moncorgé, and Knut-Andreas Lie. An implicit local time-stepping method based on cell reordering for multiphase flow in porous media. Journal of Computational Physics: X, 6:??, March 2020. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055220300032.

Lindner:2023:NSN

[LÖR23] Andreas Lindner, Baris Ölmez, and Hartmut Ruhl. Numerical simulations of the nonlinear quantum vacuum in the Heisenberg– Euler weak-field expansion. Journal of Computational Physics: X, 17:??, November 2023. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055223000021.

Lecoanet:2019:TCS

[LVB^{+}19] Daniel Lecoanet, Geoffrey M. Vasil, Keaton J. Burns, Benjamin P. Brown, and Jeffrey S. Oishi. Tensor calculus in spherical coordinates using Jacobi polynomials. Part II: Implementation and examples. Journal of Computational Physics: X, 3:??, June 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300289.

Lattanzi:2019:HLB

[LYH19] Aaron M. Lattanzi, Xiaolong Yin, and Christine M. Hrenya. A hybrid lattice boltzmann–random walk method for heat transfer in gas–solids systems. Journal of Computational Physics: X, 1:??, January 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S259005521930023X.

Maric:2021:IVF

[Mar21] Tomislav Marić. Iterative volume-of-fluid interface positioning in general polyhedrons with Consecutive Cubic Spline interpolation. Journal of Computational Physics: X, 11:??, June 2021. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S259005522100010X.

Muralikrishnan:2021:SGB

[MCF^{+}21] Sriramkrishnan Muralikrishnan, Antoine J. Cerfon, Matthias Frey, Lee F. Ricketson, and Andreas Adelmann. Sparse grid-based adaptive noise reduction strategy for particle-in-cell schemes. Journal of Computational Physics: X, 11:??, June 2021. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055221000111.

MacDonald:2020:MMM

[MMR20] Craig S. MacDonald, John A. Mackenzie, and Alison Ramage. A moving mesh method for modelling defects in nematic liquid crystals. Journal of Computational Physics: X, 8:??, September 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300172.

Morikawa:2023:CAI

[MTA23] Daniel Shigueo Morikawa, Kumpei Tsuji, and Mitsuteru Asai. Corrected ALE-ISPH with novel Neumann boundary condition and density-based particle shifting technique. Journal of Computational Physics: X, 17:??, November 2023. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055223000033.

Nagaraj:2019:DMA

[NGP^{+}19] Sriram Nagaraj, Jacob Grosek, Socratis Petrides, Leszek F. Demkowicz, and Jaime Mora. A 3D DPG Maxwell approach to nonlinear Raman gain in fiber laser amplifiers. Journal of Computational Physics: X, 2:??, March 2019. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055219300010.

Potter:2023:FHL

[PBSM23] Samuel F. Potter, Stefano Bertone, Norbert Schörghofer, and Erwan Mazarico. Fast hierarchical low-rank view factor matrices for thermal irradiance on planetary surfaces. Journal of Computational Physics: X, 17:??, November 2023. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055223000082.

Patel:2019:TDC

[PDF19] Ravi G. Patel, Olivier Desjardins, and Rodney O. Fox. Three-dimensional conditional hyperbolic quadrature method of moments. Journal of Computational Physics: X, 1: ??, January 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300058.

Parkinson:2020:MBA

[PMWK20] James R. G. Parkinson, Daniel F. Martin, Andrew J. Wells, and Richard F. Katz. Modelling binary alloy solidification with adaptive mesh refinement. Journal of Computational Physics: X, 5:??, January 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300599.

Peest:2023:MCR

[PRB23] Christian Peest, Oliver Reich, and Lena Bressel. Monte Carlo radiative transfer peel off mechanism for spatially extended detectors. Journal of Computational Physics: X, 17: ??, November 2023. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055223000100.

Ramgraber:2023:ETSa

[RBMM23a] Maximilian Ramgraber, Ricardo Baptista, Dennis McLaughlin, and Youssef Marzouk. Ensemble transport smoothing. Part I: Unified framework. Journal of Computational Physics: X, 17:??, November 2023. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055223000124.

Ramgraber:2023:ETSb

[RBMM23b] Maximilian Ramgraber, Ricardo Baptista, Dennis McLaughlin, and Youssef Marzouk. Ensemble transport smoothing. Part II: Nonlinear updates. Journal of Computational Physics: X, 17:??, November 2023. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055223000112.

Rocha:2021:FCS

[RKvdM21] I. B. C. M. Rocha, P. Kerfriden, and F. P. van der Meer. On-the-fly construction of surrogate constitutive models for concurrent multiscale mechanical analysis through probabilistic machine learning. Journal of Computational Physics: X, 9:??, January 2021. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300354.

Rahmouni:2020:MBF

[RMPA20] Lyes Rahmouni, Adrien Merlini, Axelle Pillain, and Francesco P. Andriulli. On the modeling of brain fibers in the EEG forward problem via a new family of wire integral equations. Journal of Computational Physics: X, 5:??, January 2020. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055219300642.

Silvestri:2019:FGM

[SP19] S. Silvestri and R. Pecnik. A fast GPU Monte Carlo radiative heat transfer implementation for coupling with direct numerical simulation. Journal of Computational Physics: X, 3:??, June 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300484.

Shohet:2020:HMM

[SPH^{+}20] Gil Shohet, Jacob Price, Jeffrey Haack, Mathieu Marciante, and Michael S. Murillo. Heterogeneous multiscale method for high energy-density matter: Connecting kinetic theory and molecular dynamics. Journal of Computational Physics: X, 8:??, September 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300226.

Schmidt:2019:LMR

[SPNSB19] Michael J. Schmidt, Stephen D. Pankavich, Alexis Navarre-Sitchler, and David A. Benson. A Lagrangian method for reactive transport with solid/aqueous chemical phase interaction. Journal of Computational Physics: X, 2:??, March 2019. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S259005521930037X.

Tao:2019:SGD

[TGC19] Zhanjing Tao, Wei Guo, and Yingda Cheng. Sparse grid discontinuous Galerkin methods for the Vlasov–Maxwell system. Journal of Computational Physics: X, 3:??, June 2019. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055219300381.

Tretiak:2019:AOT

[TR19] Krasymyr Tretiak and Daniel Ruprecht. An arbitrary order time-stepping algorithm for tracking particles in inhomogeneous magnetic fields. Journal of Computational Physics: X, 4:??, September 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300526.

Tsoutsanis:2019:SSA

[Tso19] Panagiotis Tsoutsanis. Stencil selection algorithms for WENO schemes on unstructured meshes. Journal of Computational Physics: X, 4:??, September 2019. CODEN JCPXAL. ISSN

2590-0552. URL http://www.sciencedirect.com/science/
article/pii/S2590055219300538.

Ullah:2019:NSM

[UkM19] Mohammed Aman Ullah and De kang Mao. Numerically simulating the missing physical dissipations in a conservative front-tracking method. Journal of Computational Physics: X, 2:??, March 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300253.

Vidal-Codina:2019:MCG

[VCSSNP19] F. Vidal-Codina, J. Saà-Seoane, N.-C. Nguyen, and J. Peraire. A multiscale continuous Galerkin method for stochastic simulation and robust design of photonic crystals. Journal of Computational Physics: X, 2:??, March 2019. CODEN JCPXAL. ISSN

2590-0552. URL http://www.sciencedirect.com/science/
article/pii/S2590055219300320.

Vasil:2019:TCS

[VLB^{+}19] Geoffrey M. Vasil, Daniel Lecoanet, Keaton J. Burns, Jeffrey S. Oishi, and Benjamin P. Brown. Tensor calculus in spherical coordinates using Jacobi polynomials. Part I: Mathematical analysis and derivations. Journal of Computational Physics: X, 3:??, June 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300290.

vanLith:2019:AFS

[vLtI19] Bart S. van Lith, Jan H. M. ten Thije Boonkkamp, and Wilbert L. IJzerman. Active flux schemes on moving meshes with applications to geometric optics. Journal of Computational Physics: X, 3:??, June 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300460.

Vittoz:2019:CWC

[VOdLL19] L. Vittoz, G. Oger, M. de Leffe, and D. Le Touzé. Comparisons of weakly-compressible and truly incompressible approaches for viscous flow into a high-order Cartesian-grid finite volume framework. Journal of Computational Physics: X, 1:??, January 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300319.

Wittwer:2023:CMS

[WA23] Lucas D. Wittwer and Sebastian Aland. A computational model of self-organized shape dynamics of active surfaces in fluids. Journal of Computational Physics: X, 17:??, November 2023. CODEN

JCPXAL. ISSN 2590-0552. URL http://www.sciencedirect.
com/science/article/pii/S2590055223000045.

Weishaupt:2019:ECF

[WJNH19] Kilian Weishaupt, Vahid Joekar-Niasar, and Rainer Helmig. An efficient coupling of free flow and porous media flow using the porenetwork modeling approach. Journal of Computational Physics: X, 1:??, January 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300277.

Wang:2019:BFE

[WL19] Xin Wang and Yingjie Liu. Back and forth error compensation and correction method for linear hyperbolic systems with application to the Maxwell’s equations. Journal of Computational Physics: X, 1:??, January 2019. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055219300307.

Wu:2020:TDL

[WPIJ20] Chuanren Wu, Ioannis Gr. Pagonakis, Stefan Illy, and John Jelonnek. Two-dimensional leapfrog scheme for trajectories of relativistic charged particles in static axisymmetric electric and magnetic field. Journal of Computational Physics: X, 8: ??, September 2020. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055220300317.

Wen:2021:IDU

[WWGS21] Xin Wen, Lian-Ping Wang, Zhaoli Guo, and Jie Shen. An improved discrete unified gas kinetic scheme for simulating compressible natural convection flows. Journal of Computational Physics: X, 11:??, June 2021. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055221000056.

Xu:2021:VCS

[XP21] YiQin Xu and Yulia T. Peet. Verification and convergence study of a spectral-element numerical methodology for fluidstructure interaction. Journal of Computational Physics: X, 10:??, March 2021. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S2590055221000019.

Zhang:2023:DGS

[ZP23] Fengrui Zhang and Yulia T. Peet. Discontinuous Galerkin spectral element method for shock capturing with summation by parts properties. Journal of Computational Physics: X, 17: ??, November 2023. CODEN JCPXAL. ISSN 2590-0552.

URL http://www.sciencedirect.com/science/article/pii/
S259005522300001X.